OpenStudy (anonymous):

Please Help Me.. Add 2 3/5 + -3 1/2 a) -6 1/10 b) -9/10 c) -4 1/10 d) -1 1/10

OpenStudy (anonymous):

@mathmate help?

OpenStudy (mathmate):

\( 2\frac{3}{4} + -3\frac{1}{2} \\ = \frac{11}{4}-\frac{7}{2} \)

OpenStudy (mathmate):

The next step would be to find a common denominator and work on the numerators.

OpenStudy (anonymous):

ok

OpenStudy (mathmate):

Oh, I just noticed that none of the answers is correct, so please check the question (or the answers).

OpenStudy (anonymous):

They have to be! and yea I know its -3/4 but idk

OpenStudy (mathmate):

Your answer is correct, guess the book/computer is wrong.

OpenStudy (anonymous):

its not.. omg this is hard

OpenStudy (anonymous):

@mathmate I found out the problem I put 3/4 instead of 3/5 ok I edited it up on the question now help me pls?

OpenStudy (mathmate):

Now that's a lot better! \( 2\frac{3}{5}+−3\frac{1}{2}\\ =\frac{13}{5}−\frac{7}{2} = ? \)

OpenStudy (anonymous):

Ok hold on @mathmate

OpenStudy (anonymous):

-0.9 and tht as a fraction is..? @mathmate

OpenStudy (anonymous):

-9/10 ? @mathmate

OpenStudy (mathmate):

Dear Juliette: thou shall not use a calculator to do fractions, because they can round off incorrectly or give recurring decimals which will make your answers incorrect. You are learning to work with fractions, so do use pencil and paper! However, you are lucky in this case because -0.9 converts well into -9/10 (which you should know how to do...lol)

OpenStudy (anonymous):

I did use a calculator and Ok tY?

OpenStudy (anonymous):

didnt*

OpenStudy (mathmate):

While you're learning fractions, put your calculator in your school bag. That's what I tell my students!

OpenStudy (mathmate):

Whew, that's good!

OpenStudy (anonymous):

@mathmate ur a teacher?

OpenStudy (mathmate):

Yes.

OpenStudy (anonymous):

Oh..

OpenStudy (mathmate):

We do all kinds of things in our lives.